# Time crystalline solitons and their stochastic dynamics in a driven-dissipative $\phi^{4}$ model

### v1

@@ Abstract
Periodical driven systems provide unique opportunities to investigate the dynamics of topological structures far from equilibrium. In this paper, we report a time-crystalline soliton (TCS) state in a driven-dissipative $\phi^{4}$ model. This state exhibits spontaneous breaking of discrete time-translational symmetry alongside spatial soliton behavior. During time evolution, the soliton pattern periodically oscillates between kink and anti-kink configurations. We further study TCS dynamics under noise, demonstrating that soliton random walk can induce a dynamical transition between two distinct $Z_{2}$ symmetry-breaking time-crystalline phases in time domain. Finally, we examine the annihilation of two spatially separated TCSs under noise. Importantly, in contrast to the confined behavior of time-crystalline monopoles reported in [Phys. Rev. Lett. 131, 056502 (2023)], the dynamics of time-crystalline solitons is deconfined despite the nonequilibrium nature of our model: the statistically averaged annihilation time scales as a power law with the solitons’ initial separation.
@@ I Introduction
Beyond identifying new phases, a profound fundamental question arises: What are the elementary excitations residing above such far-from-equilibrium, symmetry-broken states? In equilibrium condensed matter, elementary excitations define the low-energy physics – phonons propagate through crystals, while topological excitations like domain walls (solitons) mediate transitions between degenerate ground states in Ising ferromagnets breaking discrete $Z_{2}$ symmetry. Extending this concept to highly excited, dissipative nonequilibrium phases remains a significant challenge and opportunity. Recent work has begun exploring topological excitations within DTC frameworks. Slow parameter ramping, for instance, can induce a dynamical transition between two distinct $Z_{2}$ symmetry-breaking DTC phases, interpreted as a temporal analog of a spatial soliton[10]. Furthermore, studies of monopole excitations in a time-crystalline spin ice revealed dynamical confinement – a direct consequence of the underlying nonequilibrium drive, contrasting sharply with the deconfined dynamics of their equilibrium counterparts[35].
@@ II Model and method
Where $\eta$ is the dissipation strangth fixed as $\eta=0.5$ throughout this paper. $\xi(x,t)$ is a zero-mean $\langle\xi(x,t)\rangle_{\xi}=0]$ random field representing thermal fluctuations. And the stochastic variables satisfy
@@ III.1 Time-crystalline solitons
We start the time evolution with a topological soliton initial state, i.e. $\phi(x,0)=\sqrt{A}\mathrm{tanh}[\frac{\sqrt{A}(x-x_{0})}{2}]$ under open boundary condition. We note that the long-time dynamics does not depend on the specific value of $A$. (See Appendix B for comparison with different initial states) . Then we monitor the evolution of the scalar field $\phi(x,t)$ under the periodic driving $\epsilon=J\mathrm{cos}(\omega_{0}t)$ according to Equation 2. Figure 1 (a) is a plot of $\phi(x,t)$ in the two-dimensional spacetime, where the field configuration periodically oscillates in time. And the soliton core (the bright central line) preserves its localized structure during the time evolution. In Figure1 (b), we can directly observe that $\phi(x_{0},t)$ exhibits a persistent oscillation with a period of $2T$, i.e. twice that of the driving period ($T=2\pi/\omega_{0}$). And the Fourier spectrum with dominant peak at $\omega_{0}/2$ confirms the periodic doubling dynamics. This periodic doubling phenomenon is a manifestation of spontaneous $Z_{2}$ discrete time translational symmetry breaking[$\phi(x,t)=\phi(x,t+2T)\neq\phi(x,t+T)$]. Figure1(c) depicts snapshot of $\phi(x,t)$ at time slice $t_{0}$ and $t_{0}+T$. The soliton profile exhibits kink-antikink oscillation during one period of driving and it returns to its original profile only after two periods of driving , i.e., $2T$, which actually reflects the SSB of DTTS. In summary, TCS is a non-equilibrium topological localized structure with DTC order.
Notably, the TCS acts as a dynamical domain wall separating two degenerate DTC phases. Each DTC phase exhibits robust $2T$-periodicity while being related to each other by shifting a driving period $T$ along the temporal direction. This temporal degeneracy emerges from $Z_{2}$ symmetry breaking of DTTS. Figure1(d) dipicts time evolution of $\phi(x,t)$ at two symmetric positions ($x_{0}^{\prime}=-x_{0}$ ). It can be seen $\phi(x_{0},t)$ and $\phi(x_{0}^{\prime},t)$ can be connected to each other by shifting a driving period $T$, which can be considered as two $Z_{2}$ symmetry breaking “degenerate” DTC phases. Crucially, we demonstrate below that stochastic motion of the TCS dynamically mediates transitions between these two degenerate DTC phases, revealing a novel noise-driven topological mechanism for temporal phase switching.
@@
III.2 Stochastic dynamics of TCS and instanton-like excitations
@@
We now investigate the stochastic dynamics of the time-crystalline soliton under noise perturbations. Figure 3(a) displays the long-time spatiotemporal evolution of the TCS under a single noise trajectory, revealing clear stochastic motion of the TCS center driven by thermal fluctuations. Quantitative analysis of the noise-induced diffusion is presented in Figure3(b), where the ensemble-averaged mean squared displacement follows $\langle x(t)^{2}\rangle_{\xi}\sim t$, confirming normal diffusive behavior. The statistical ensemble average is taken over $2000$ independent noise trajectories. Figure 3(c) provides a magnified view of the temporal evolution from (a), while Figure3(d) compares the field evolution $\phi(x_{0},t)$ along two distinct paths indicated in (c). Crucially, the blue curve (crossing the soliton center) exhibits a $\pi$-phase shift, whereas the red curve (avoiding the core) shows no such phase inversion. This $\pi$-phase shift serves as a dynamical transition between the two $Z_{2}$ symmetry-breaking degenerate DTC phases, resembling an instanton excitation in quantum field theory[40] and thereby constituting a topological defect in the temporal domain. In Figure3(b), we observe that the TCS exhibits significantly reduced diffusion compared to equilibrium solitons in the undriven system. This suppression of diffusion arises because the activation of a DTC instanton requires additional energy.
@@ III.3 Deconfinement of two spatially separated TCSs
In summary, by investigating a driven-dissipative $\phi^{4}$ model, we have uncovered a exotic time-crystalline soliton (TCS) state, which is a non-equilibrium topological localized structure that simultaneously breaks discrete time translational symmetry and exhibits spatial solitonic structure. The TCS manifests as a dynamical domain wall separating two degenerate $Z_{2}$ symmetry-breaking DTC phases, with its profile periodically oscillating between kink and antikink configurations during one driving period. Crucially, the stochastic motion of the TCS induces dynamical transitions between these two degenerate phases, realizing an instanton-like excitation in the time domain. Furthermore, despite the strongly non-equilibrium nature of the system, the TCSs exhibit deconfined dynamics—the annihilation time of soliton-antisoliton pairs follows a power-law scaling with their initial separation, in sharp contrast to the exponentially confined dynamics observed in 2D non-equilibrium spin ice systems.
@@ Appendix A The algorithm for solving the stochastic Langevin equation
This appendix details the numerical algorithm employed to solve the stochastic Landau-Ginzburg equation 2 . The second-order time derivative is first decomposed into two first-order equations:
@@
Then we discretize spacetime as $x_{i}=i\Delta x$, $t_{l}=l\Delta t$, where $i,l=0,1,2,\cdot\cdot\cdot$.
$\phi_{i}^{l}=\phi(x_{i},t_{l}),$ (6)
$y_{i}^{l}=y(x_{i},t_{l}),$ (7)
A predictor-corrector scheme is adopted to ensure convergence and handle the stochastic term robustly. The update procedure for one full time step is as follows:
Step 1: Predictor Step for the Field
The field is predicted at the next time step using a simple Euler update:
@@
Spatial derivatives are computed using the standard central difference scheme on both the current and predicted configurations:
@@
Step 2: Update of the Auxiliary Field $y(x,t)$
Equation (5) is discretized using an implicit-explicit approach for stability, averaging the spatial derivatives and treating dissipation implicitly:
@@
where is $g_{i}^{l}$ a random number satisfying the Gaussian distribution with $\mathcal{N}(0,1)$: $\langle g_{i}^{l}\rangle_{\bm{\xi}}=0$, $\langle(g_{i}^{l})^{2}\rangle_{\bm{\xi}}=1$.
Step 3: Corrector Step for the Field
The field configuration is finally updated using a trapezoidal rule, which constitutes the corrector step of the algorithm:
@@
This numerical scheme combines the stability of an implicit treatment of dissipation with the accuracy of a predictor-corrector method for the stochastic dynamics, ensuring robust integration of the stochastic partial differential equation.
@@
Appendix B CONVERGENCE OF NUMERICAL RESULTS
B.1 Finite $\Delta t$ and finite $\Delta x$

### v2

@@ Abstract
Periodically driven systems provide unique opportunities to investigate the dynamics of topological excitations far from equilibrium. In this paper, we report a time-crystalline soliton (TCS) state in a driven-dissipative $\phi^{4}$ model. This state exhibits spontaneous breaking of discrete time-translational symmetry while simultaneously displaying spatial soliton behavior. During time evolution, the soliton pattern periodically oscillates between kink and anti-kink configurations. We further study TCS dynamics under noise, demonstrating that soliton random walk can induce a dynamical transition between two distinct $Z_{2}$ symmetry-breaking time-crystalline phases in time domain. Finally, we examine the annihilation of two spatially separated TCSs under noise. Importantly, in contrast to the confined behavior of time-crystalline monopoles reported in [Phys. Rev. Lett. 131, 056502 (2023)], the dynamics of time-crystalline solitons is deconfined despite the nonequilibrium nature of our model: the statistically averaged annihilation time scales as a power law with the solitons’ initial separation.
@@ I Introduction
Beyond identifying new phases, a profound fundamental question arises: What are the elementary excitations residing above such far-from-equilibrium, symmetry-broken states? In equilibrium condensed matter, elementary excitation is closely related to spontaneous symmetry breaking. Elementary excitations define the low-energy physics – phonons propagate through crystals, while topological excitations like domain walls (solitons) mediate transitions between degenerate ground states in Ising ferromagnets breaking discrete $Z_{2}$ symmetry. Extending this concept to highly excited, dissipative nonequilibrium phases remains a significant challenge and opportunity. Recent work has begun exploring topological excitations within DTC frameworks. Slow parameter ramping, for instance, can induce a dynamical transition between two distinct $Z_{2}$ symmetry-breaking DTC phases, interpreted as a temporal analog of a spatial soliton[10]. Furthermore, studies of monopole excitations in a time-crystalline spin ice revealed dynamical confinement – a direct consequence of the underlying nonequilibrium drive, contrasting sharply with the deconfined dynamics of their equilibrium counterparts[35].
@@ II Model and method
Where $\eta$ is the dissipation strength fixed as $\eta=0.5$ throughout this paper. $\xi(x,t)$ is a zero-mean $\langle\xi(x,t)\rangle_{\xi}=0]$ random field representing thermal fluctuations. And the stochastic variables satisfy
@@ III.1 Time-crystalline solitons
We start the time evolution with a topological soliton initial state, i.e. $\phi(x,0)=\sqrt{A}\mathrm{tanh}[\frac{\sqrt{A}(x-x_{0})}{2}]$ under open boundary condition. We note that the long-time dynamics does not depend on the specific value of $A$. (See Appendix C for comparison with different initial states) . Then we monitor the evolution of the scalar field $\phi(x,t)$ under the periodic driving $\epsilon=J\mathrm{cos}(\omega_{0}t)$ according to Equation 2. Figure 1 (a) is a plot of $\phi(x,t)$ in the two-dimensional spacetime, where the field configuration periodically oscillates in time. And the soliton core (the bright central line) preserves its localized structure during the time evolution. In Figure1 (b), we can directly observe that $\phi(x_{0},t)$ exhibits a persistent oscillation with a period of $2T$, i.e. twice that of the driving period ($T=2\pi/\omega_{0}$). And the Fourier spectrum with dominant peak at $\omega_{0}/2$ confirms the periodic doubling dynamics. This periodic doubling phenomenon is a manifestation of spontaneous $Z_{2}$ discrete time translational symmetry breaking[$\phi(x,t)=\phi(x,t+2T)\neq\phi(x,t+T)$]. Figure1(c) depicts snapshot of $\phi(x,t)$ at time slice $t_{0}$ and $t_{0}+T$. The soliton profile exhibits kink-antikink oscillation during one period of driving and it returns to its original profile only after two periods of driving , i.e., $2T$, which actually reflects the SSB of DTTS. In summary, TCS is a non-equilibrium topological localized structure with DTC order.
Notably, the TCS acts as a dynamical domain wall separating two degenerate DTC phases. Each DTC phase exhibits robust $2T$-periodicity while being related to each other by shifting a driving period $T$ along the temporal direction. This temporal degeneracy emerges from $Z_{2}$ symmetry breaking of DTTS. Figure1(d) dipicts time evolution of $\phi(x,t)$ at two symmetric positions ($x_{0}^{\prime}=-x_{0}$ ). It can be seen $\phi(x_{0},t)$ and $\phi(x_{0}^{\prime},t)$ can be connected to each other by shifting a driving period $T$, which can be considered as two $Z_{2}$ symmetry breaking “degenerate” DTC phases. We demonstrate below that stochastic motion of the TCS dynamically mediates transitions between these two degenerate DTC phases in the temporal direction.
@@
III.2 Stochastic dynamics of TCS and its diffusive behavior
@@
We now investigate the stochastic dynamics of the time-crystalline soliton under noise perturbations. Figure 3(a) displays the long-time spatiotemporal evolution of the TCS under a single noise trajectory, revealing clear stochastic motion of the TCS center driven by thermal fluctuations. Quantitative analysis of the noise-induced diffusion is presented in Figure3(b), where the ensemble-averaged mean squared displacement follows $\langle x(t)^{2}\rangle_{\xi}\sim t$, confirming normal diffusive behavior. The statistical ensemble average is taken over $2000$ independent noise trajectories. Figure 3(c) provides a magnified view of the temporal evolution from (a), while Figure3(d) compares the field evolution $\phi(x_{0},t)$ along two distinct paths indicated in (c). Crucially, the blue curve (crossing the soliton center) exhibits a $\pi$-phase shift, whereas the red curve (avoiding the core) shows no such phase inversion. This $\pi$-phase shift serves as a dynamical transition between the two $Z_{2}$ symmetry-breaking degenerate DTC phases, and thereby constituting a topological defect in the temporal domain. In Figure3(b), we observe that the TCS exhibits significantly reduced diffusion compared to equilibrium solitons in the undriven system. This suppression of diffusion arises because the activation of a DTC defect requires additional energy.
@@ III.3 Deconfinement of two spatially separated TCSs
In summary, by investigating a driven-dissipative $\phi^{4}$ model, we have uncovered a exotic time-crystalline soliton (TCS) state, which is a non-equilibrium topological localized structure that simultaneously breaks discrete time translational symmetry and exhibits spatial solitonic structure. The TCS manifests as a dynamical domain wall separating two degenerate $Z_{2}$ symmetry-breaking DTC phases, with its profile periodically oscillating between kink and antikink configurations during one driving period. Crucially, the stochastic motion of the TCS induces dynamical transitions between these two degenerate phases, realizing a topological excitation in the time domain. Furthermore, despite the strongly non-equilibrium nature of the system, the TCSs exhibit deconfined dynamics—the annihilation time of soliton-antisoliton pairs follows a power-law scaling with their initial separation, in sharp contrast to the exponentially confined dynamics observed in 2D non-equilibrium spin ice systems.
@@ Appendix A The algorithm for solving the stochastic Langevin equation
The stochastic Landau-Ginzburg equation 2 can be solved by the standard Heun method. In every time step, we need to update the field configuration $\phi(x,t)$. First, we decompose the second-order time derivative into two first-order equations:
@@
Then we discretize spacetime as $x_{i}=i\Delta x$, $t_{l}=l\Delta t$, where $i,l=0,1,2,\cdot\cdot\cdot$. Then the field configurations can be discretized as $\phi_{i}^{l}=\phi(x_{i},t_{l})$, $y_{i}^{l}=y(x_{i},t_{l})$.
We calculate an intermediate configuration $\widetilde{\phi}_{i}^{l+1}$ as
@@
Spatial derivatives are computed using the standard central difference scheme on both the current and intermediate configurations:
@@
Equation (5) is discretized as:
@@
where $g_{i}^{l}$ is a random number satisfying the Gaussian distribution with $\mathcal{N}(0,1)$: $\langle g_{i}^{l}\rangle_{\bm{\xi}}=0$, $\langle(g_{i}^{l})^{2}\rangle_{\bm{\xi}}=1$.
In the final step, we calculate the field configuration at $(l+1)$th time step,
@@
Appendix B $J$-$\omega_{0}$ phase diagram
In this section, we explore the range of parameters $J$ and $\omega_{0}$ for which the DTC order occurs. In summary, we plot the dynamical phase diagram in terms of the driving amplitude $J$ and the driving frequency $\omega_{0}$ in Fig.6, from which we can find that small driving amplitude favors a synchronized trivial phase, while large driving amplitude leads to an DTC phase. We also notice that low driving frequency leads to more unstable chaotic behavior of the system. The DTC order can be found for sufficiently large and fast driving. Even though in the maintext, we focused only on a specific point in the parameter space $(J=5,\omega_{0}=\pi)$, the proposed TCS is a robust phase that exists in a large regime of the phase diagram. This observation makes it easier to realize the TCS in a more realistic experimental setup.
Figure 6: Zero temperature dynamical phase diagram in terms of the driving amplitude $J$ and the driving frequency $\omega_{0}$.
Appendix C CONVERGENCE OF NUMERICAL RESULTS
C.1 Finite $\Delta t$ and finite $\Delta x$